# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ IV Training methodologies
This section comprehensively overviews state‑of‑the‑art approaches to improve LLMs and advance their reasoning capabilities. Specifically, the section covers various techniques—from general training methods and architectural innovations (including various Mixture-of-Experts and attention modifications) to reinforcement learning strategies and retrieval‑augmented generation methods. It further details advanced chain‑of‑thought approaches, test‑time compute scaling strategies, and other innovative techniques that address challenges such as self‑improvement, efficient inference, and data deduplication. These techniques are instrumental in advancing LLM reasoning by enabling models to process better, synthesize, and apply knowledge in complex, multi-step tasks.
@@
TableIV presents a comprehensive overview of training methodologies for LLMs, categorizing them into groups that span general training approaches, mixture-of-experts (MoE) and architectural innovations, retrieval-augmented generation (RAG), chain-of-thought and self-improvement techniques, as well as test-time compute scaling and distillation methods. Notably, many of these approaches have been developed and reported between 2024 and 2025, reflecting a period of rapid innovation. By detailing the key techniques and their main impacts, this table provides valuable insights into how contemporary training paradigms are evolving to improve model performance, efficiency, and adaptability across diverse tasks.
@@
Table V focuses specifically on reinforcement learning (RL) methods applied to LLMs, showcasing a range of techniques from direct preference optimization to complex self-correction and constrained RL strategies. Emphasizing contributions from 2024 to 2025, the table highlights how RL-based approaches have been leveraged to enhance reasoning accuracy, reduce training time, and optimize overall model behavior. Collectively, the methodologies outlined in this table underscore the critical role of reinforcement learning in advancing LLM capabilities and bridging the performance gap with human-like reasoning.
IV-A Agent Q
@@
Putta et al. [64] presents a novel framework for empowering large language models to perform complex, multi-step reasoning in dynamic, interactive environments. Recognizing the limitations of traditional supervised pre-training and expert demonstration fine-tuning—which often lead to compounding errors and limited exploration—the proposed method combines guided Monte Carlo Tree Search (MCTS) with a self-critique mechanism and an off-policy Direct Preference Optimization (DPO) algorithm. This integrated approach, dubbed Agent Q, enables LLM agents to learn effectively from successful and unsuccessful trajectories, significantly improving task performance. Validated in a simulated e-commerce environment (WebShop), Agent Q, when applied to a Llama-3 70B model, boosts zero-shot success rates from 18.6% to 81.7%, further reaching 95.4% with online search and even surpasses average human performance. These results underscore the potential of this framework to revolutionize autonomous agent decision-making, with applicability extending to various agentic workflows such as code execution and API interactions.
@@
The Agent Q produces a sequence of composite actions $\{a_{t}\}$ by considering its history $h_{t}$ at each time step. It builds upon a ReAct-style agent [105] but incorporates a dedicated planning step (PlanReact) [106] along with additional elements. The agent’s behavior is decomposed into four main components:
@@
IV-B Low-Rank Adaptation (LoRA)
Biderman et al. [65] comprehensively compare Low-Rank Adaptation (LoRA) and full finetuning for large language models in programming and mathematics. By conducting experiments on Llama-2 7B and 13B models—utilizing both instruction tuning (with roughly 100K prompt-response pairs) and continued pretraining (using approximately 10B unstructured tokens)—the study demonstrates that while LoRA, which fine-tunes only low-rank perturbations to selected weight matrices, acts as a strong regularizer by maintaining diverse generations and preserving performance on tasks outside the target domain, it substantially underperforms full finetuning on the target domain tasks such as those measured by HumanEval and GSM8K. The results further reveal that full finetuning learns perturbations with a rank that is 10 to 100 times greater than typical LoRA configurations, which may explain its superior accuracy and sample efficiency in specialized tasks.
@@ IV-B Low-Rank Adaptation (LoRA)
IV-C Cognitive Architectures for Language Agents (CoALA)
@@
Sumers et al. [66] introduced Cognitive Architectures for Language Agents (CoALA), a structured framework that integrates insights from cognitive science and symbolic AI with contemporary large language models to enhance the capabilities of language agents. CoALA posits that by endowing language agents with modular memory systems—including working, long-term, and procedural memory—and a clearly defined action space for both internal (e.g., reasoning and memory updates) and external (e.g., tool or API interactions) operations, these agents can more effectively manage grounding and reasoning tasks. The framework further incorporates a generalized decision-making process that involves proposing, evaluating, and selecting actions, ensuring that the agent’s behavior aligns with its objectives. This structured approach organizes a broad array of recent developments in language agents and lays out actionable directions for future research, ultimately contextualizing modern language agents within the rich tradition of cognitive architectures and marking a promising path toward more versatile, general intelligence in language-based AI systems.
IV-D Learn-by-interact
@@
Learn-by-interact [67] is a novel data-centric framework designed to enhance the adaptability of large language model (LLM) agents across diverse digital environments without the need for human annotations. The framework synthesizes realistic agent-environment interaction trajectories by leveraging available documentation—such as manuals, API references, and tutorials. A key innovation in this process is ”backward construction,” which involves summarizing or abstracting interaction histories to generate clear, aligned instructions from observed behaviors. The authors validate the effectiveness of this synthetic data in both fine-tuning and in-context learning settings, achieving notable improvements: up to 12.2% enhancement for in-context learning with Claude-3.5 and 19.5% for fine-tuning with Codestral-22B. Extensive experiments conducted on benchmarks like SWE-bench, WebArena, OSWorld, and Spider2-V across coding, web, and desktop domains underscore the robustness of the approach. Ablation studies further confirm the critical contribution of backward construction and the superiority of their retrieval pipeline over traditional methods such as retrieval-augmented generation. Overall, Learn-by-interact offers a scalable solution for synthesizing high-quality agent data, paving the way for more capable and adaptive UI agents in real-world applications.
IV-E Supervised Fine-tuning (SFT)
@@
Chu et al. [68] presented a systematic study comparing supervised fine-tuning (SFT) and reinforcement learning (RL) in terms of their impact on model generalization and memorization, particularly within both textual rule-based and visual domains. Using the arithmetic reasoning card game GeneralPoints and the real-world navigation environment V-IRL as testbeds, the study evaluates the performance of a Llama-3.2-Vision-11B model under different training regimes: SFT-only, RL-only, and a hybrid SFT→RL pipeline. The results reveal that while SFT tends to memorize the training data—leading to strong in-distribution performance but poor out-of-distribution generalization—RL, mainly when guided by outcome-based rewards, effectively generalizes across both rule-based textual and visual tasks. Notably, RL enhances generalization and improves the model’s visual recognition capabilities. However, the findings also emphasize that SFT remains a critical precursor to RL, as it stabilizes the model’s output format and ensures that the backbone model adheres to instructions, enabling RL to realize its performance gains fully. Additional experiments with multiple verification iterations (or ”Reject Sampling”) further demonstrate that this hybrid approach can boost generalization by up to approximately 6%.
IV-F Mixture of Experts (MoE)
@@
MoE [69] comprises multiple ”expert” neural networks alongside a router that probabilistically directs incoming tokens to the most appropriate expert. Unlike traditional models where every parameter is active for each inference, MoE leverages its router to selectively activate a subset of experts, enabling the model to encompass many parameters while keeping inference computationally efficient. This selective activation is governed by carefully managed load balancing, ensuring that all experts contribute effectively during training. The approach is versatile, extending beyond language models to vision applications, as exemplified by Mixtral 8x7B, which incorporates a staggering 46.7 billion parameters but only engages 12.8 billion during inference. This design paradigm underscores the potential of MoE to scale model capacity significantly without proportionally increasing the computational burden during deployment [42, 48, 107, 108].
@@
MoE layers are typically distributed across multiple GPUs or devices to handle large numbers of experts. Techniques such as Expert Parallelism (EP) split experts among different devices. In a Transformer model, each MoE layer is placed within a feed-forward block (or a sub-block, depending on the architecture), and the gating network determines how tokens are routed to the experts. Frameworks like GShard implement Top2 gating and use Switch or SwiGLU operations to process each token’s expert assignments in a distributed fashion. This significantly allows for efficient training and inference when scaling to large models [109, 110].
IV-F1 Heterogeneous Mixture of Experts (HMoE)
@@
Wang et al. [70] presents a novel approach to the Mixture of Experts (MoE) paradigm by introducing a Heterogeneous Mixture of Experts (HMoE) framework, which departs from traditional homogeneous designs by employing experts of varying sizes. The key insight is that differing token complexities in natural language processing tasks benefit from experts with specialized capacities rather than a one-size-fits-all model. To this end, the authors propose a unique training objective—leveraging a penalty mechanism that preferentially activates smaller experts—to improve computational efficiency and parameter utilization. Experimental evaluations reveal that HMoE not only achieves a lower loss but also requires fewer activated parameters than conventional MoE models, demonstrating superior performance across a range of pre-training benchmarks. This innovative method underscores the potential for more adaptive expert allocation strategies in large-scale language modeling.
IV-F2 Expert-Specialized Fine-Tuning (ESFT)
@@
Tang et al. [71] investigate parameter-efficient fine-tuning (PEFT) methods for LLMs with sparse architectures, explicitly focusing on Mixture-of-Experts (MoE) models. While most PEFT research has centered on dense-architecture models, this work explores the unique challenges and opportunities in fine-tuning MoE models. The authors first analyze the dispersion of activated experts across tasks, revealing that the routing distribution is highly concentrated for individual tasks but varies significantly across different tasks. Building on this observation, they introduce Expert-Specialized Fine-Tuning (ESFT), a method that selectively tunes the most relevant experts for a given downstream task while freezing the others. Experimental results demonstrate that ESFT not only improves tuning efficiency—reducing memory usage by up to 90% (from 28.6GB to as low as 2.57GB for Token and 3.20GB for Gate ESFT) and training time by up to 30% (from 28.5 minutes to 19.8 minutes)—but also achieves performance comparable to or even surpassing full-parameter fine-tuning and outperforms LoRA by up to 10%. Moreover, the study highlights that MoE models with finer-grained experts are better at selecting task-relevant combinations, whereas the approach is less practical for coarse-grained MoEs like Mixtral. These findings underscore the benefits of task-specific expert specialization in PEFT for MoE-based LLMs and provide valuable insights into optimizing resource usage and training efficiency.
IV-F3 Mixture-of-Agents
@@
Wang et al. [72] propose a novel Mixture-of-Agents (MoA) approach to harness the collective strengths of multiple LLMs for enhanced natural language generation and understanding. In the MoA framework, a layered architecture is employed where each layer consists of several LLM agents that take the aggregated outputs from the previous layer as input, thereby iteratively refining the final response. This method leverages different models’ diversity and complementary capabilities, enabling them to generate more accurate and contextually rich outputs. Empirical evaluations on benchmarks such as AlpacaEval 2.0, MT-Bench, and FLASK demonstrate that MoA models achieve state-of-the-art performance—for instance, a score of 65.1% on AlpacaEval 2.0 compared to 57.5% by GPT-4 Omni. Furthermore, a variant called MoA-Lite outperforms GPT-4 Turbo by 4% while being twice as cost-efficient, although the layered aggregation process can lead to increased latency in generating the first token.
IV-F4 Multimodal Mixture of Experts
@@
Li et al. [73] introduce Uni-MoE, a unified multimodal large language model that leverages a sparse Mixture-of-Experts (MoE) architecture to efficiently process a wide array of modalities—including audio, speech, image, text, and video—within a single framework. Unlike traditional MoE approaches that employ homogeneous experts and are limited in modality coverage, Uni-MoE utilizes modality-specific encoders with connectors to build a unified multimodal representation, enabling both modality-level data parallelism and expert-level model parallelism. The proposed model is trained using a progressive, three-phase strategy: first, cross-modality alignment is achieved via various connectors; second, modality-specific experts are trained with cross-modality instruction data to activate expert preferences; and finally, the system is fine-tuned using Low-Rank Adaptation (LoRA) on mixed multimodal instruction data. Extensive evaluations across ten vision and audio tasks demonstrate that Uni-MoE matches or outperforms existing unified multimodal models, significantly reducing performance bias in handling diverse and mixed datasets.
IV-G Multihead Latent Attention (MLA)

### v2

@@ VI Training methodologies
This section comprehensively overviews state‑of‑the‑art approaches to improve LLMs and advance their reasoning capabilities. Specifically, the section covers various techniques from general training methods and architectural innovations (including various Mixture-of-Experts and attention modifications) to reinforcement learning strategies and retrieval‑augmented generation methods. It further details advanced chain‑of‑thought approaches, test‑time compute scaling strategies, and other innovative techniques that address challenges such as self‑improvement, efficient inference, and data deduplication. These techniques are instrumental in advancing LLM reasoning by enabling models to process better, synthesize, and apply knowledge in complex, multi-step tasks.
@@
Table. IV presents a comprehensive overview of training methodologies for LLMs, categorizing them into groups that span general training approaches, mixture-of-experts (MoE) and architectural innovations, retrieval-augmented generation (RAG), chain-of-thought and self-improvement techniques, as well as test-time compute scaling and distillation methods. Notably, many of these approaches have been developed and reported between 2024 and 2025, reflecting a period of rapid innovation. By detailing the key techniques and their main impacts, this table provides valuable insights into how contemporary training paradigms are evolving to improve model performance, efficiency, and adaptability across diverse tasks.
@@
Table. V focuses specifically on reinforcement learning (RL) methods applied to LLMs, showcasing a range of techniques from direct preference optimization to complex self-correction and constrained RL strategies. Emphasizing contributions from 2024 to 2025, the table highlights how RL-based approaches have been leveraged to enhance reasoning accuracy, reduce training time, and optimize overall model behavior. Collectively, the methodologies outlined in this table underscore the critical role of reinforcement learning in advancing LLM capabilities and bridging the performance gap with human-like reasoning.
VI-A Agent Q
@@
Putta et al. [putta2024agent] presents a novel framework for empowering large language models to perform complex, multi-step reasoning in dynamic, interactive environments. Recognizing the limitations of traditional supervised pre-training and expert demonstration fine-tuning which often lead to compounding errors and limited exploration the proposed method combines guided Monte Carlo Tree Search (MCTS) with a self-critique mechanism and an off-policy Direct Preference Optimization (DPO) algorithm. This integrated approach, dubbed Agent Q, enables LLM agents to learn effectively from successful and unsuccessful trajectories, significantly improving task performance. Validated in a simulated e-commerce environment (WebShop), Agent Q, when applied to a Llama-3 70B model, boosts zero-shot success rates from 18.6% to 81.7% , further reaching 95.4% with online search and even surpassing average human performance. These results underscore the potential of this framework to revolutionize autonomous agent decision-making, with applicability extending to various agentic workflows such as code execution and API interactions.
@@
The Agent Q produces a sequence of composite actions $\{a_{t}\}$ by considering its history $h_{t}$ at each time step. It builds upon a ReAct-style agent [yao2023react] but incorporates a dedicated planning step (PlanReact) [liu2023bolaa] along with additional elements. The agent’s behavior is decomposed into four main components:
@@
VI-B Low-Rank Adaptation (LoRA)
Biderman et al. [biderman2024lora] comprehensively compare Low-Rank Adaptation (LoRA) and full finetuning for large language models in programming and mathematics. By conducting experiments on Llama-2 7B and 13B models utilizing both instruction tuning (with roughly 100K prompt-response pairs) and continued pretraining (using approximately 10B unstructured tokens) the study demonstrates that while LoRA, which fine-tunes only low-rank perturbations to selected weight matrices, acts as a strong regularizer by maintaining diverse generations and preserving performance on tasks outside the target domain, it substantially underperforms full finetuning on the target domain tasks such as those measured by HumanEval and GSM8K. The results further reveal that full finetuning learns perturbations with a rank 10 to 100 times greater than typical LoRA configurations, which may explain its superior accuracy and sample efficiency in specialized tasks.
@@ VI-B Low-Rank Adaptation (LoRA)
VI-C Cognitive Architectures for Language Agents (CoALA)
@@
Sumers et al. [sumers2023cognitive] introduced Cognitive Architectures for Language Agents (CoALA), a structured framework that integrates insights from cognitive science and symbolic AI with contemporary large language models to enhance the capabilities of language agents. CoALA posits that by endowing language agents with modular memory systems including working, long-term, and procedural memory and a clearly defined action space for both internal (e.g., reasoning and memory updates) and external (e.g., tool or API interactions) operations, these agents can more effectively manage grounding and reasoning tasks. The framework further incorporates a generalized decision-making process that involves proposing, evaluating, and selecting actions, ensuring that the agent’s behavior aligns with its objectives. This structured approach organizes a broad array of recent developments in language agents and lays out actionable directions for future research, ultimately contextualizing modern language agents within the rich tradition of cognitive architectures and marking a promising path toward more versatile, general intelligence in language-based AI systems.
VI-D Learn-by-interact
@@
Learn-by-interact [su2025learn] is a novel data-centric framework designed to enhance the adaptability of large language model (LLM) agents across diverse digital environments without the need for human annotations. The framework synthesizes realistic agent-environment interaction trajectories by leveraging documentation such as manuals, API references, and tutorials. A key innovation in this process is ”backward construction,” which involves summarizing or abstracting interaction histories to generate clear, aligned instructions from observed behaviors. The authors validate the effectiveness of this synthetic data in both fine-tuning and in-context learning settings, achieving notable improvements: up to 12.2% enhancement for in-context learning with Claude-3.5 and 19.5% for fine-tuning with Codestral-22B. Extensive experiments conducted on benchmarks like SWE-bench, WebArena, OSWorld, and Spider2-V across coding, web, and desktop domains underscore the robustness of the approach. Ablation studies further confirm the critical contribution of backward construction and the superiority of their retrieval pipeline over traditional methods such as retrieval-augmented generation. Overall, Learn-by-interact offers a scalable solution for synthesizing high-quality agent data, paving the way for more capable and adaptive UI agents in real-world applications.
VI-E Supervised Fine-tuning (SFT)
@@
Chu et al. [chu2025sft] presented a systematic study comparing supervised fine-tuning (SFT) and reinforcement learning (RL) in terms of their impact on model generalization and memorization, particularly within both textual rule-based and visual domains. Using the arithmetic reasoning card game GeneralPoints and the real-world navigation environment V-IRL as testbeds, the study evaluates the performance of a Llama-3.2-Vision-11B model under different training regimes: SFT-only, RL-only, and a hybrid SFT→RL pipeline. The results reveal that while SFT tends to memorize the training data, leading to strong in-distribution performance, but poor out-of-distribution generalization, RL, mainly when guided by outcome-based rewards, effectively generalizes across both rule-based textual and visual tasks. Notably, RL enhances generalization and improves the model’s visual recognition capabilities. However, the findings also emphasize that SFT remains a critical precursor to RL, as it stabilizes the model’s output format and ensures that the backbone model adheres to instructions, enabling RL to realize its performance gains fully. Additional experiments with multiple verification iterations (or ”Reject Sampling”) further demonstrate that this hybrid approach can boost generalization by up to approximately 6%.
VI-F Mixture of Experts (MoE)
@@
MoE [jiang2024mixtral] comprises multiple ”expert” neural networks alongside a router that probabilistically directs incoming tokens to the most appropriate expert. Unlike traditional models, where every parameter is active for each inference, MoE leverages its router to selectively activate a subset of experts, enabling the model to encompass many parameters while keeping inference computationally efficient. This selective activation is governed by carefully managed load balancing, ensuring that all experts contribute effectively during training. The approach is versatile, extending beyond language models to vision applications, as exemplified by Mixtral 8x7B, which incorporates a staggering 46.7 billion parameters but only engages 12.8 billion during inference. This design paradigm underscores the potential of MoE to scale model capacity significantly without proportionally increasing the computational burden during deployment [deepseekai2025, liu2024deepseek, puigcerver2023sparse, riquelme2021scaling].
@@
MoE layers are typically distributed across multiple GPUs or devices to handle large numbers of experts. Techniques such as Expert Parallelism (EP) split experts among different devices. In a Transformer model, each MoE layer is placed within a feed-forward block (or a sub-block, depending on the architecture), and the gating network determines how tokens are routed to the experts. Frameworks like GShard implement Top2 gating and use Switch or SwiGLU operations to process each token’s expert assignments in a distributed fashion. This significantly allows for efficient training and inference when scaling to large models [lepikhin2020gshard, shazeer2017outrageously].
VI-F1 Heterogeneous Mixture of Experts (HMoE)
@@
Wang et al. [wang2024hmoe] presents a novel approach to the Mixture of Experts (MoE) paradigm by introducing a Heterogeneous Mixture of Experts (HMoE) framework, which departs from traditional homogeneous designs by employing experts of varying sizes. The key insight is that differing token complexities in natural language processing tasks benefit from experts with specialized capacities rather than a one-size-fits-all model. To this end, the authors propose a unique training objective leveraging a penalty mechanism that preferentially activates smaller experts to improve computational efficiency and parameter utilization. Experimental evaluations reveal that HMoE not only achieves a lower loss but also requires fewer activated parameters than conventional MoE models, demonstrating superior performance across a range of pre-training benchmarks. This innovative method underscores the potential for more adaptive expert allocation strategies in large-scale language modeling.
VI-F2 Expert-Specialized Fine-Tuning (ESFT)
@@
Tang et al. [tang2024understanding] investigate parameter-efficient fine-tuning (PEFT) methods for LLMs with sparse architectures, explicitly focusing on Mixture-of-Experts (MoE) models. While most PEFT research has centered on dense-architecture models, this work explores the unique challenges and opportunities in fine-tuning MoE models. The authors first analyze the dispersion of activated experts across tasks, revealing that the routing distribution is highly concentrated for individual tasks but varies significantly across different tasks. Building on this observation, they introduce Expert-Specialized Fine-Tuning (ESFT), a method that selectively tunes the most relevant experts for a given downstream task while freezing the others. Experimental results demonstrate that ESFT not only improves tuning efficiency reducing memory usage by up to 90% (from 28.6GB to as low as 2.57GB for Token and 3.20GB for Gate ESFT) and training time by up to 30% (from 28.5 minutes to 19.8 minutes) but also achieves performance comparable to or even surpassing full-parameter fine-tuning and outperforms LoRA by up to 10%. Moreover, the study highlights that MoE models with finer-grained experts are better at selecting task-relevant combinations, whereas the approach is less practical for coarse-grained MoEs like Mixtral. These findings underscore the benefits of task-specific expert specialization in PEFT for MoE-based LLMs and provide valuable insights into optimizing resource usage and training efficiency.
VI-F3 Mixture-of-Agents
@@
Wang et al. [wang2024mixture] propose a novel Mixture-of-Agents (MoA) approach to harness the collective strengths of multiple LLMs for enhanced natural language generation and understanding. In the MoA framework, a layered architecture is employed where each layer consists of several LLM agents that take the aggregated outputs from the previous layer as input, thereby iteratively refining the final response. This method leverages different models’ diversity and complementary capabilities, enabling them to generate more accurate and contextually rich outputs. Empirical evaluations on benchmarks such as AlpacaEval 2.0, MT-Bench, and FLASK demonstrate that MoA models achieve state-of-the-art performance for instance, a score of 65.1% on AlpacaEval 2.0 compared to 57.5% by GPT-4 Omni. Furthermore, a variant called MoA-Lite outperforms GPT-4 Turbo by 4% while being twice as cost-efficient, although the layered aggregation process can lead to increased latency in generating the first token.
VI-F4 Multimodal Mixture of Experts
@@
Li et al. [li2025uni] introduce Uni-MoE, a unified multimodal large language model that leverages a sparse Mixture-of-Experts (MoE) architecture to efficiently process a wide array of modalities, including audio, speech, image, text, and video, within a single framework. Unlike traditional MoE approaches that employ homogeneous experts and are limited in modality coverage, Uni-MoE utilizes modality-specific encoders with connectors to build a unified multimodal representation, enabling both modality-level data parallelism and expert-level model parallelism. The proposed model is trained using a progressive, three-phase strategy: first, cross-modality alignment is achieved via various connectors; second, modality-specific experts are trained with cross-modality instruction data to activate expert preferences; and finally, the system is fine-tuned using Low-Rank Adaptation (LoRA) on mixed multimodal instruction data. Extensive evaluations across ten vision and audio tasks demonstrate that Uni-MoE matches or outperforms existing unified multimodal models, significantly reducing performance bias in handling diverse and mixed datasets.
VI-G Multihead Latent Attention (MLA)